OpenStudy (anonymous):

QUESTION #1- what is the geometric mean of 5 and 15? QUESTION #2- 12 is the geometric mean of 9 and what number?

OpenStudy (mertsj):

gm of ab = sqrt(ab)

OpenStudy (anonymous):

1 sqrt (5*15) = sqrt 75 2. 12^2 = 144 144 / 9 = 16 12 is geometric mean of 9 and 16